Pat's present age is three-fourths of Ken's present age. In 15 years the sum of their ages will be 58, Find their present ages.
Answer by gonzo(654)

let x = pat's present age.
let y = ken's present age.
---
Pat's present age is three-fourths of Ken's present age.
x = 3/4 * y
---
In 15 years the sum of their ages will be 58
x + 15 + y + 15 = 58
combine like terms to get:
x + y + 30 = 58
subtract 30 from both sides of equation to get:
x + y = 28
---
substitute 3/4 * y for x to get:
3/4 * y + y = 28
multiply both sides of equation by 4 to get:
3*y + 4*y = 4*28
combine like terms to get:
7*y = 4*28
divide both sides of equation by 7 to get:
y = 4*28/7
simplify to get:
y = 16
---
since x = 3/4*y, then:
x = (3/4)*16 = 3*16/4 = 3*4 = 12
---
you have:
x = 12
y = 16
---
pat is 12
ken is 16
pat's age is 3/4 * ken's age.
in 15 years, pat will be 27 and ken will be 31.
sum of their ages will be 58
everything checks out.
answer is:
pat is 12
ken is 16
---